# When Deep Learning Meets Polyhedral Theory: A Survey

### v3

@@ 4.3.3 A class of intermediate formulations
The previous sections observed that the big-$M$ formulation (4) is compact, but may offer a weak convex relaxation, while the extended formulation (5) offers the tightest possible convex relaxation for an individual unit, at the expense of a much larger formulation. Kronqvist et al. (2022, 2021) present a strategy for obtaining formulations intermediate to (4) and (5) in terms of both size and strength. The key idea is to partition ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}$ into a number of aggregated variables, ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}=\sum_{p=1}^{P}\hat{x}_{p}$. Each auxiliary variable $\hat{x}_{p}$ is defined as a sum of a subset of the $j$-th weighted inputs $\hat{x}_{p}=\sum_{j\in\mathbb{S}_{p}}w_{i,j}^{l}h_{j}^{l-1}$, with $\mathbb{S}_{1},...,\mathbb{S}_{P}$ partitioning $\{1,...,n_{l-1}\}$. This technique can be applied to the ReLU function, giving the convex hull over the directions defined by $\hat{x}_{p}$ (Tsay et al., 2021):
@@ 4.3.4 Cutting plane methods: Trading variables for inequalities
$\breve{L}^{l-1}_{j}$ =$=\begin{cases}L^{l-1}_{j}&w^{l}_{i,j}\geq 0\\ U^{l-1}_{j}&w^{l}_{i,j}<0\end{cases}\quad\mathrm{and}\quad\breve{U}^{l-1}_{j}=\begin{cases}U^{l-1}_{j}&w^{l}_{i,j}\geq 0\\ L^{l-1}_{j}&w^{l}_{i,j}<0\end{cases}$
@@ 4.4.1 Projecting the big-$M$ and ideal MILP formulations
(Projecting the big-$M$). Note that the LP relaxation given by (4a4d) maintains the variables $z^{l}_{i}$ in the formulation, though they are no longer required to satisfy integrality. Since these variables are “auxiliary” and are no longer necessary to encode the nonconvexity of the problem, they can be projected out without altering the quality of the convex relaxation. Doing this yields what is commonly known as the “triangle” or “$\Delta$” relaxation (Salman et al., 2019):
@@ 4.4.2 Dual decomposition methods
Note that the final three constraints apply the big-$M$/triangle relaxation (8) to each ReLU activation function. The dual problem can then be solved via subgradient-based methods, proximal algorithms, or, more recently, a projected gradient descent method applied to a nonconvex reformulation of the problem (Bunel et al., 2020c).
@@ 4.4.3 Fourier-Motzkin elimination and propagation algorithms
Figure 10: Convex approximations for the ReLU function commonly used by propagation algorithms, given as a function of the preactivation function $\hat{h_{i}^{l}}$. The ReLU applies $h_{i}^{l}=\max(0,\hat{h_{i}^{l}})$.
@@
Such tools use methods such as abstract transformers to propagate polyhedral bounds, i.e., zonotopes, through the layers of a neural network. DeepZ (Singh et al., 2018), Fast-Lin (Weng et al., 2018), and Neurify (Wang et al., 2018a) employ a parallel approximation, with the latter also implementing a branch-and-bound procedure towards completeness. Subsequently, DeepPoly (Singh et al., 2019b) and CROWN (Zhang et al., 2018a) select between the zero and identity approximations by minimizing over-approximation area. OSIP (Hashemi et al., 2021) selects between the three approximations using optimization: approximations for a layer are select jointly to minimise bounds for the following layer. These technologies are also compatible with interval bounds, propagating box domains (Mirman et al., 2018). Interestingly, bounds on neural network weights can also be propagated using similar methods, allowing reachability analysis of Bayesian neural networks (Wicker et al., 2020).
@@
s.t. $\mathcal{L}_{k}({\bm{h}}^{l-2})\leq h^{l-1}_{k}\leq\mathcal{G}_{k}({\bm{h}}^{l-2})\forall k\in\{1,...,n_{l-1}\}.$
@@ 4.5.2 Extending to other activation functions
While this survey focuses on neural networks with piecewise linear activation functions, we note that recent research has also studied smooth activation functions with a similar aim. For example, optimization over smooth activation functions can be handled by piecewise linear approximation and conversion to MILP (Sildir and Aydin, 2022). Researchers have also studied convex/concave bounds for nonlinear activation functions, which can then be embedded in spatial branch-and-bound procedures (Schweidtmann and Mitsos, 2019, Wilhelm et al., 2022). In contrast to MILP formulations for ReLU neural networks, these problems are typically nonlinear programs that must be solved via spatial branch and bound.
Propagation methods (Singh et al., 2018, Zhang et al., 2018a) can also naturally handle general activation functions: given convex polytopic bounds for an activation function, these tools can propagate them through network layers using the same techniques. For example, Fastened CROWN (Lyu et al., 2020) employs a set of search heuristics to quickly select linear upper and lower bounds on ReLU, sigmoid, and hyperbolic tangent activation functions. More recent work combines linear bounds with tailored branching rules for networks with general nonlinear functions (Shi et al., 2024). Tighter polyhedral bounds can be employed, such as piecewise linear upper and lower bounds (Benussi et al., 2022).
@@ 4.5.3 Extending to adversarial training
As described in Section 1, the training of neural networks seeks to minimise a measure of distance between the output $y$ and the correct output $\hat{y}$. For instance, if this distance is prescribed as a loss function $\mathcal{L}(y,\hat{y})$, this corresponds to solving the training optimization problem:
$\underset{\{{\bm{W}}^{l}\}_{l\in{\mathbb{L}}},\{{\bm{b}}^{l}\}_{l\in{\mathbb{L}}}}{\mathrm{min}}\mathcal{L}(y,\hat{y}).$ (10)
Further details about the training problem and solution methods are described in Section 5. An interesting direction here is to produce networks with bounded global (Cisse et al., 2017, Leino et al., 2021) or local (Huang et al., 2021) Lipschitz constants—a paradigm with some connections to polyhedral theory. Nevertheless, given the focus of this section on optimization, we aim to briefly outline how verification techniques can be embedded in training. Specifically, solutions or bounds to the verification problem (Section 4.1.1) provide a metric of how robust a trained neural network is to perturbations. These metrics can be embedded in the training problem to obtain a more robust network during training, often resulting in a bilevel training problem. For instance, the verification problem (3) can be embedded as a lower-level problem, giving the robust optimization problem:
$\underset{\{{\bm{W}}^{l}\}_{l\in{\mathbb{L}}},\{{\bm{b}}^{l}\}_{l\in{\mathbb{L}}}}{\mathrm{min}}\quad\underset{||x-\hat{x}||\leq\epsilon}{\mathrm{max}}\quad\mathcal{L}(y=f(x),\hat{y}).$
Solving these problems generally involves either bilevel optimization, or computing an adversarial solution/bound at each training step, conceptually similar to the robust cutting plane approach. Madry et al. (2018) proposed this formulation and solved the nonconvex inner problem using gradient descent, thereby losing a formal certification of robustness. These approaches may also benefit from reformulation strategies, such as by taking the dual of the inner problem and using any feasible solution as a bound (Wong and Kolter, 2018). The resulting models are not only more robust, but several works have also found it to be empirically easier to verify robustness in them (Mirman et al., 2018, Wong and Kolter, 2018).
@@ 5.1.1 Problem setting and solution scheme
Note that when a neural network $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}f$ has only one hidden layer, its behavior is almost completely determined by the sign of each component of the vector ${\bm{W}}^{1}x-{\bm{b}}^{1}.$ These are the cases of ReLU activations $\sigma(z)=\max\{0,z\}$ and LTU activations $\sigma(z)=\text{sgn}(z)$. The training algorithms we show here heavily exploit this observation and construct $({\bm{W}}^{1},{\bm{b}}^{1})$ by embedding in this phase a hyperplane partition problem based on the sample $(\tilde{{\bm{x}}}_{i},\tilde{{\bm{y}}}_{i})_{i=1}^{D}$. While the focus of this survey is mainly devoted to ReLU activations, we also discuss some selected cases with LTU activations as they share some similar ideas.
@@ 5.1.2 LTU activations and variable number of nodes
  • Thanks to the universal approximation theorem (Hornik et al., 1989), with enough nodes in the hidden layer, we can always obtain a neural network $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}f$ with zero classification error; although, this can lead to over-fitting.
@@ 5.2 Convex reformulations in regularized training problems
Using that $P\leq 2r(e(D-1)/r)^{r}$, where $r=\mbox{rank}(\tilde{{\bm{X}}})$, the authors note that the formulation (19) yields a training algorithm with complexity $O({n_{0}}^{3}r^{3}(D/r)^{3r})$. Note that if one fixes $r$, the resulting algorithm runs in polynomial time. In particular, fixing ${n_{0}}$ fixes the rank of $\tilde{{\bm{X}}}$ and results in a polynomial time algorithm as well. In contrast, the algorithm by Arora et al. (2018) discussed in Section 5.1.3 remains exponential even after fixing ${n_{0}}$. Moreover, Froese and Hertrich (2024) showed that the training problem is NP-Hard even for fixed ${n_{0}}$. This apparent contradiction is explained by two key components of the convex reformulation: the regularization term and the presence of a “large enough” number of hidden neurons. This facilitates the exponential improvement of the training algorithm with respect to Arora et al. (2018).
@@ 5.4 Polyhedral encoding of multiple training problems
As in the previous section, we omit the biases ${\bm{b}}$ to simplify notation, as all parameters can be included in ${\bm{W}}$. Let us assume the class of neural networks $F$ in (11) are restricted to have bounded parameters (we assume they lie in the interval $[-1,1]$), and let us assume the sample has been normalized in such a way that $(\tilde{{\bm{x}}}_{i},\tilde{{\bm{y}}}_{i})\in[-1,1]^{{n_{0}}+{n_{L+1}}}$. Furthermore, let $N$ be the dimension of $\Theta$ (the number of parameters in the neural network). With this notation, we define the following.
@@
over the domain $[-1,1]^{n_{0}}\times\Theta\times[-1,1]^{n_{L+1}}$.
@@
This result is very abstract in nature but possesses some interesting features. Firstly, it encodes (approximately) every possible training problem arising from data in $[-1,1]^{{n_{0}}+{n_{L+1}}}$ using a benign dependency on $D$: the polytope size depends only linearly on $D$, while a discretized enumeration of all the possible samples of size $D$ would be exponential in $D$. Secondly, every possible ERM problem appears in a face of the polytope; this suggests a strong geometric structure across different ERM problems. Lastly, this result is applicable to a wide variety of network architectures; in order to obtain an architecture-specific result, it suffices to compute the corresponding value of $\mathcal{K}$ and plug it in. Regarding this last point, the authors computed the constant $\mathcal{K}$ for various well-known architectures and obtained the results of Table 4.

### v4

@@ 4.3.3 A class of intermediate formulations
The previous sections observed that the big-$M$ formulation (4) is compact, but may offer a weak convex relaxation, while the extended formulation (5) offers the tightest possible convex relaxation for an individual unit, at the expense of a much larger formulation. Kronqvist et al. (2025, 2021) present a strategy for obtaining formulations intermediate to (4) and (5) in terms of both size and strength. The key idea is to partition ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}$ into a number of aggregated variables, ${\bm{w}}_{i}^{l}{\bm{h}}^{l-1}=\sum_{p=1}^{P}\hat{x}_{p}$. Each auxiliary variable $\hat{x}_{p}$ is defined as a sum of a subset of the $j$-th weighted inputs $\hat{x}_{p}=\sum_{j\in\mathbb{S}_{p}}w_{i,j}^{l}h_{j}^{l-1}$, with $\mathbb{S}_{1},...,\mathbb{S}_{P}$ partitioning $\{1,...,n_{l-1}\}$. This technique can be applied to the ReLU function, giving the convex hull over the directions defined by $\hat{x}_{p}$ (Tsay et al., 2021):
@@ 4.3.4 Cutting plane methods: Trading variables for inequalities
$\breve{L}^{l-1}_{j}$ $=\begin{cases}L^{l-1}_{j}&w^{l}_{i,j}\geq 0\\ U^{l-1}_{j}&w^{l}_{i,j}<0\end{cases}\quad\mathrm{and}\quad\breve{U}^{l-1}_{j}=\begin{cases}U^{l-1}_{j}&w^{l}_{i,j}\geq 0\\ L^{l-1}_{j}&w^{l}_{i,j}<0\end{cases}$
@@ 4.4.1 Projecting the big-$M$ and ideal MILP formulations
(Projecting the big-$M$). Note that the LP relaxation given by (4a4d) maintains the variables $z^{l}_{i}$ in the formulation, though they are no longer required to satisfy integrality. Since these variables are “auxiliary” and are no longer necessary to encode the nonconvexity of the problem, they can be projected out without altering the quality of the convex relaxation. Doing this yields what is commonly known as the “triangle” or $\Delta$” relaxation (Salman et al., 2019):
@@ 4.4.2 Dual decomposition methods
Note that the final three constraints apply the big-MM/triangle relaxation (8) to each ReLU activation function. The dual problem can then be solved via subgradient-based methods, proximal algorithms, or, more recently, a projected gradient descent method applied to a nonconvex reformulation of the problem (Bunel et al., 2020c).
@@ 4.4.3 Fourier-Motzkin elimination and propagation algorithms
[…]
@@
Such tools use methods such as abstract transformers to propagate polyhedral bounds, i.e., zonotopes, through the layers of a neural network. DeepZ (Singh et al., 2018), Fast-Lin (Weng et al., 2018), and Neurify (Wang et al., 2018a) employ a parallel approximation, with the latter also implementing a branch-and-bound procedure towards completeness. Subsequently, DeepPoly (Singh et al., 2019b) and CROWN (Zhang et al., 2018a) select between the zero and identity approximations by minimizing over-approximation area. OSIP (Hashemi et al., 2021) selects between the three approximations using optimization: approximations for a layer are select jointly to minimise bounds for the following layer. These technologies are also compatible with interval bounds, propagating box domains (Mirman et al., 2018). Interestingly, bounds on neural network weights can also be propagated using similar methods, allowing reachability analysis of Bayesian neural networks (Wicker et al., 2020) and reachability analysis of the weights themselves during training steps (Sosnin et al., 2025, Wicker et al., 2025).
@@
s.t. $\mathcal{L}_{k}({\bm{h}}^{l-2})\leq h^{l-1}_{k}\leq\mathcal{G}_{k}({\bm{h}}^{l-2})\forall k\in\{1,.,n_{l-1}\}.$
@@ 4.5.2 Extending to other activation functions
While this survey focuses on neural networks with piecewise linear activation functions, we note that recent research has also studied smooth activation functions with a similar aim. For example, optimization over smooth activation functions can be handled by piecewise linear approximation and conversion to MILP (Sildir and Aydin, 2022). Researchers have also studied convex/concave bounds for nonlinear activation functions, which can then be embedded in spatial branch-and-bound procedures (Schweidtmann and Mitsos, 2019, Wilhelm et al., 2022). Notably, Carrasco and Muñoz (2024) take a similar approach to the extended formulations described in Section 4.3 to derive tight formulations for convex and S-shaped activation functions; the relaxations were later implemented for the hyperbolic tangent activation function in a branch-and-bound solver by Witte et al. (2025). In contrast to MILP formulations for ReLU neural networks, these problems are typically nonlinear programs that must be solved via spatial branch and bound.
Propagation methods (Singh et al., 2018, Zhang et al., 2018a) can also naturally handle general activation functions: given convex polytopic bounds for an activation function, these tools can propagate them through network layers using the same techniques. For example, Fastened CROWN (Lyu et al., 2020) employs a set of search heuristics to quickly select linear upper and lower bounds on ReLU, sigmoid, and hyperbolic tangent activation functions. More recent work combines linear bounds with tailored branching rules for networks with general nonlinear functions (Shi et al., 2025). Tighter polyhedral bounds can be employed, such as piecewise linear upper and lower bounds (Benussi et al., 2022).
@@ 4.5.3 Extending to adversarial training
[…]
Further details about the training problem and solution methods are described in Section 5. An interesting direction here is to produce networks with bounded global (Cisse et al., 2017, Leino et al., 2021) or local (Huang et al., 2021) Lipschitz constants—a paradigm with some connections to polyhedral theory. Nevertheless, given the focus of this section on optimization, we aim to briefly outline how verification techniques can be embedded in training. Specifically, solutions or bounds to the verification problem (Section 4.1.1) provide a metric of how robust a trained neural network is to perturbations. These bounds can be used directly to bound the reachable set of network weights during training, given some perturbations (Sosnin et al., 2025, Wicker et al., 2025). Alternatively, verification-based metrics can be embedded in the training problem to obtain a more robust network during training, often resulting in a bilevel training problem. For instance, the verification problem (3) can be embedded as a lower-level problem, giving the robust optimization problem:
[…]
Solving these problems generally involves either bilevel optimization, or computing an adversarial solution/bound at each training step, conceptually similar to the robust cutting plane approach. Madry et al. (2018) proposed this formulation and solved the nonconvex inner problem using gradient descent, thereby losing a formal certification of robustness. These approaches may also benefit from reformulation strategies, such as by taking the dual of the inner problem and using any feasible solution as a bound (Wong and Kolter, 2018) or by using projection-based approaches (Francobaldi and Lombardi, 2025). The resulting models are not only more robust, but several works have also found it to be empirically easier to verify robustness in them (Mirman et al., 2018, Wong and Kolter, 2018).
@@ 5.1.1 Problem setting and solution scheme
Note that when a neural network $f$ has only one hidden layer, its behavior is almost completely determined by the sign of each component of the vector ${\bm{W}}^{1}x-{\bm{b}}^{1}.$ These are the cases of ReLU activations $\sigma(z)=\max\{0,z\}$ and LTU activations $\sigma(z)=\text{sgn}(z)$. The training algorithms we show here heavily exploit this observation and construct $({\bm{W}}^{1},{\bm{b}}^{1})$ by embedding in this phase a hyperplane partition problem based on the sample $(\tilde{{\bm{x}}}_{i},\tilde{{\bm{y}}}_{i})_{i=1}^{D}$. While the focus of this survey is mainly devoted to ReLU activations, we also discuss some selected cases with LTU activations as they share some similar ideas.
@@ 5.1.2 LTU activations and variable number of nodes
  • Thanks to the universal approximation theorem (Hornik et al., 1989), with enough nodes in the hidden layer, we can always obtain a neural network $f$ with zero classification error; although, this can lead to over-fitting.
@@ 5.2 Convex reformulations in regularized training problems
[…]
@@ 5.4 Polyhedral encoding of multiple training problems
As in the previous section, we omit the biases $\bm{b}$} to simplify notation, as all parameters can be included in $\bm{W}$}. Let us assume the class of neural networks $F$ in (11) are restricted to have bounded parameters (we assume they lie in the interval [1,1][-1,1]), and let us assume the sample has been normalized in such a way that $(\tilde{{\bm{x}}}_{i},\tilde{{\bm{y}}}_{i})\in[-1,1]^{{n_{0}}+{n_{L+1}}}$. Furthermore, let $N$ be the dimension of $\Theta$ (the number of parameters in the neural network). With this notation, we define the following.
@@
over the domain $1,1]^{n_{0}}\times\Theta\times[-1,1]^{n_{L+1}}$.
@@
This result is very abstract in nature but possesses some interesting features. Firstly, it encodes (approximately) every possible training problem arising from data in $1,1]^{{n_{0}}+{n_{L+1}}}$ using a benign dependency on $D$: the polytope size depends only linearly on $D$, while a discretized enumeration of all the possible samples of size $D$ would be exponential in $D$. Secondly, every possible ERM problem appears in a face of the polytope; this suggests a strong geometric structure across different ERM problems. Lastly, this result is applicable to a wide variety of network architectures; in order to obtain an architecture-specific result, it suffices to compute the corresponding value of $\mathcal{K}$ and plug it in. Regarding this last point, the authors computed the constant $\mathcal{K}$ for various well-known architectures and obtained the results of Table 4.